OpenStudy (anonymous):

Jacob and Roman are standing next to each other. Jacob, who is 180cm tall, has a shadow that is 30cm long. If Roman is 168cm tall, approximately how long is his shadow? 6cm 18cm 28cm 42cm

OpenStudy (anonymous):

45 cm / 180 cm = x / 152 cm x = 45 * 152 / 180 = 38 cm

OpenStudy (anonymous):

Oh I thought it would be 28cm O.o

OpenStudy (anonymous):

(: No

OpenStudy (anonymous):

But on my quiz it doesnt have that answer ....

OpenStudy (anonymous):

It Should Be 38 o.O

OpenStudy (anonymous):

Roman's shadow should be 38cm?

OpenStudy (anonymous):

Yes